# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ 2.3 Uniqueness Results for Non-rotating Bodies
Here we provide a proof outline. One can refer to [13, Lemma 12] to understand some step(s) omitted here. We first note that the one-to-one correspondence between a central density of zero and a total mass of zero is trivial, as both correspond to the vacuum (zero function). Hence we can now turn to the case where mass is positive.
The idea of Lieb and Yau is: let $G=(0,\infty)$, $D=\{\alpha\mid\sigma^{\alpha}\text{ is a minimizer with total mass }m\in G\}$. For each $m\in G$, we know by Theorem 2.6 and Proposition 2.43 there is a unique minimizer $\rho_{m}$, and a unique central density $\alpha_{m}$ (see also Corollary 2.44). We let $\Gamma:G\rightarrow D$ denote this map from $m\in G$ to $\alpha_{m}\in D$.
Then in [13, Lemma 12], one can show $\Gamma$ (or $\Gamma^{-1}$) is a homeomorphism between $G$ and $D$, and $D$ turns out to be closed in $\mathbb{R}^{+}$. $G$ is connected, so is $D$, hence $D$ is a closed interval in $\mathbb{R}^{+}$. Since $\Gamma^{-1}$ is a homeomorphism, then $D$ can only be $\mathbb{R}^{+}$. ∎
@@ 3 Scaling Relations between Stars with Different Mass
The goal of this section is to provide a description of scaling method and scaling relations between solutions with different mass, which can help us to find more quantitative properties. We have been using those results in the previous arguments and will continue to use them in the following. We assume the polytropic equations of state $P(\rho)=K\rho^{\gamma}$ with index $\gamma>\frac{4}{3}$ holds true from this section.
@@ 3.1 Scaling relations between solutions to partial differential equations
where $\pi$ is the pressure and $v>0$ is the viscosity coefficient. For simplicity we assume $v=1$ is constant. Formally, we take the divergence of the first equation and use the second equation to eliminate some terms and obtain $\Delta\pi=-\operatorname{div}((u\cdot\nabla)u)$, and then ${div}((u\cdot\nabla)u)+\operatorname{div}(\nabla\pi)=0$. Then we can introduce Leray projection operator $\mathbb{P}=Id+\nabla(-\Delta)^{-1}\operatorname{div}$ [8] to get rid of the pressure term from the first equation and obtain:
@@
However, when we consider the reduced Euler-Poisson equations (EP’) with $\omega=0$, or the corresponding Euler-Lagrange equations (EL) with $J=0$ (non-rotating case), since the pressure $P(\rho)=K\rho^{\gamma}$ or the function $A^{\prime}(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ is not linear and there is an extra term $V_{\rho},c$ may not be chosen arbitrarily. With an abuse of notation, we assume $g_{A,B}(x)=Ag(Bx)$. Let a function $\sigma(x)$ be a solution to (EL) $A^{\prime}(\sigma(x))=\left[V_{\sigma}(x)+\lambda\right]_{+}$, then we have $A^{\prime}\left(\sigma_{\widetilde{A},B}(x)\right)=\frac{K\gamma}{\gamma-1}\left(\sigma_{\widetilde{A},B}(x)\right)^{\gamma-1}=\widetilde{A}^{\gamma-1}A^{\prime}(\sigma(Bx)),V_{\sigma_{A,B}}(x)=\int\frac{\widetilde{A}\sigma(By)}{|x-y|}d^{3}y=\widetilde{A}B^{-2}\int\frac{\sigma(By)}{|Bx-By|}d^{3}(By)=\widetilde{A}B^{-2}V_{\sigma}(Bx)$. Hence if $\widetilde{A}^{\gamma-1}=\widetilde{A}B^{-2}$ or $B=\widetilde{A}^{\frac{2-\gamma}{2}}$, we will have $\sigma_{\widetilde{A},B}$ is also a solution to (EL) with $\lambda_{\widetilde{A},B}=\widetilde{A}B^{-2}:A^{\prime}\left(\sigma_{\widetilde{A},B}(x)\right)=\left[V_{\sigma_{\bar{A},B}}(x)+\lambda_{\widetilde{A},B}\right]_{+}$.
Moreover, if we know $g$ has mass $m$ (again without loss of generality we assume $m\neq 0$. unless otherwise specified) and hope $g_{A,B}$ has mass 1. Then $\int g_{A,B}=AB^{-3}$ $\int g(Bx)d^{3}(Bx)=AB^{-3}m=1$. Therefore, we can solve $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$.
@@
For the case of Euler-Lagrange equations (EL), one can see the arguments above. For the case of the reduced Euler-Poisson equations (EP’), notice (EP’) becomes $(\nabla P(g(x)))-g(x)\left(\nabla V_{g}(x)\right)=0$, and $\left(\nabla P\left(g_{A,B}(x)\right)\right)-g_{A,B}(x)\left(\nabla V_{g_{A,B}}(x)\right)=A^{\gamma}B\left(\nabla_{Bx}P(g(Bx))\right)-A^{2}B^{-1}g(Bx)\left(\nabla_{Bx}V_{g}(Bx)\right)$, thus the arguments are essentially the same. ∎
Therefore, to solve (EL), we can also assume a solution has mass 1. Notice the radial solution with certain mass is actually minimizer and unique (see Proposition 2.43 and Proposition 2.51) and vice versa. It can help us to discuss more properties and relations between minimizers with different mass in the next subsection.
@@ 3.2 Scaling relations between minimizers of $E_{0}(\rho)$ with different mass
Recall (39) that $E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}=\int A(\rho)-\frac{1}{2}\iint\frac{\rho(x)\rho(y)}{|x-y|}d^{3}xd^{3}y$, where $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$.
Let $\sigma$ be in admissible class, then we have
$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int A\left(\sigma_{\widetilde{A},B}\right)$
$=\int\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}$
$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}d^{3}(Bx)$
@@
$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}d^{3}xd^{3}y$
$=\widetilde{A}^{2}B^{-5}\int\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}d^{3}(Bx)d^{3}(By)$
@@
$\int\sigma_{\widetilde{A},B}$ $=\widetilde{A}B^{-3}\int\sigma(Bx)d^{3}(Bx)$
$=\widetilde{A}B^{-3}\int\sigma$
Assume $\sigma_{m}$ is a minimizer of $E_{0}(\rho)$ on $m{R}\left(\mathbb{R}^{3}\right)$, if $A^{\gamma}B^{-3}=A^{2}B^{-5}$ or $B=A^{\frac{2-\gamma}{2}}$, we will have $\left(\sigma_{m}\right)_{A,B}$ is a minimizer of $A^{\gamma}B^{-3}E_{0}(\rho)$ on $AB^{-3}m{R}\left(\mathbb{R}^{3}\right)$ (one can show it by contradiction arguments), and thereby a minimizer of $E_{0}(\rho)$ on $AB^{-3}m{R}\left(\mathbb{R}^{3}\right)$. If we hope $\left(\sigma_{m}\right)_{A,B}$ has mass $1$, then $AB^{-3}m=1$. Therefore, we can solve $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, which are compatible with the results in subsection 3.1 as one can expect.
@@
Thanks to Theorem 2.6, we know $\|\sigma\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ and spt $\sigma$ is contained in a ball of radius $R_{1}$, therefore, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{m}=\frac{C_{1}}{A}$ and spt $\sigma_{m}$ is contained in a ball of radius $R_{m}=BR_{1}$. Since $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, if we further assume $\gamma>2$, we know $\lim\limits_{m\rightarrow 0}A=+\infty$ and $\lim\limits_{m\rightarrow 0}B=0$, thus $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ and the size of $\sigma_{m}$’s support will go to 0 when $m\rightarrow 0$, with rates $m^{\frac{2}{3\gamma-4}}$ and $m^{\frac{\gamma-2}{3\gamma-4}}$ respectively. Notice $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus it gives the decay rate and strengthens Corollary 2.44. However, when $\gamma<2$, $\lim\limits_{m\rightarrow 0}B=\infty$, the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread). It coincides with a result mentioned in Lieb and Yau’s paper [13, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$. In their paper quantum mechanics (fermions case) is discussed.
@@
Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies $E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}$ $E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$, where $\rho(x):=\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}},\rho$ has mass 1.
@@
The result $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$-differentiable at $\rho_{m}$ comes from Lemma 2.8. Also easy to check $\rho\in W$ and has mass 1, and $E_{0}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Given $\tau_{m}\in P_{\infty}\left(\rho_{m}\right)$, one can check $\tau:=\left(\tau_{m}\right)_{A,B}\in P_{\infty}(\rho)$. By definition of variational derivatives, one has
$\int E_{0}^{\prime}\left(\rho_{m}\right)\tau_{m}$ $=\lim_{t\rightarrow\infty}\frac{E_{0}\left(\rho_{m}+t\tau_{m}\right)-E_{0}\left(\rho_{m}\right)}{t}$
@@
$=A^{-\gamma}B^{3}\int E_{0}^{\prime}(\rho)\tau$
$=A^{-\gamma-1}\int E_{0}^{\prime}(\rho)(x)\tau_{m}(Bx)d^{3}(Bx)$
$=A^{-\gamma-1}\int E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)\tau_{m}(x)d^{3}x$
Since there are enough functions $\tau_{m}$ in $P_{\infty}\left(\rho_{m}\right)$, thus we have $E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}$ $E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$. ∎
@@
Given $\rho\in W$, there is another way to show the result: we already know the expression of $E_{0}^{\prime}(\rho)=A^{\prime}(\rho(x))-V_{\rho}(x)$. Let $\rho_{m}(x)=\frac{1}{A}\rho\left(\frac{x}{B}\right),A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, one can also compute explicitly that $A^{\prime}\left(\rho_{m}\right)(x)=A^{\prime}\left(\frac{1}{A}\rho\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}A^{\prime}(\rho)\left(\frac{1}{B}x\right),V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\rho}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\rho}\left(\frac{1}{B}x\right)$, and obtain again $E_{0}^{\prime}\left(\rho_{m}\right)(x)=\frac{1}{A^{\gamma-1}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$.
Appendix A Appendix
Here we introduce some preliminary knowledge which are used in the paper.
A.1 Calculus of Variations in Vector Spaces
Since the functional we deal with is not differentiable in any usual sense, we begin with a less restrictive notion of differentiability.
Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$. Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)$, such that $u+tv\in U,0\leq t<\epsilon(v)\}$. This set is a cone in $X$. If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$. We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $\mu$ on $X(P)$ such that $\forall v\in P$
$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=\mu(v),$
The linearity of $\mu$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $\mu$ by $E_{u}^{\prime}$—with $P$ understood from context.
We say $E$ has a weak local minimum at $u\in U$ if for each $v\in P(u)$, there exists $\{t_{n}\}$ such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$ and $E\left({u+t_{n}v}\right)\geq E(u)$ for $n$ large enough. And $u$ is called weak local minimizer. We discuss more about the motivation of weak local minimizer in [7].
Let $W\subset\left\{u\in U\middle|g(u)=M\right\}$, where $g$ is a linear functional on $X$ and $M$ is a constant. Then similarly, we say $E$ has a weak local minimum at $u$ restricted to $W$ if $u\in W$ and for each $v\in P(u)$, there exists $t_{n}$} such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$, and for for $n$ large enough, we have $u+t_{n}v\in W$, and $E\left({u+t_{n}w}\right)\geq E(u)$. Furthermore, we have a result for weak local minimum restricted to $W$:
Theorem A.1 ((Generalized) Lagrange Multiplier Theorem [1, Section 2] [7, Appendix A]).
Suppose $E$ is $P$-differentiable at $u\in U$ and that its restriction to $W$ has a weak local minimum at $u$. Suppose also that $P$ is convex and that there is a $u_{0}\in P$ such that $u_{0}\in P$ and $g\left(u_{0}\right)\neq 0$. Then there is a unique constant $\lambda$ such that $E_{u}^{\prime}(v)\geq\lambda g(v)$, all $v\in P$.
A.2 Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this thesis.
Proposition A.2 (Interpolation Inequality [3, Section 4.2]).
If $f\in L^{p}\cap L^{q}$ with $1\leq p\leq q\leq\infty$, then $f\in L^{r}$ for all $r,p\leq r\leq q$, and the following interpolation inequality holds:
$\|f\|_{r}\leq\|f\|_{p}^{\alpha}\|f\|_{q}^{1-\alpha},\text{ where }\frac{1}{r}=\frac{\alpha}{p}+\frac{1-\alpha}{q},0\leq\alpha\leq 1$
Proposition A.3 (Hardy-Littlewood-Sobolev Inequality [2, Theorem 1.7]).

### v2

@@ 2.3 Uniqueness Results for Non-rotating Bodies
Here we provide a proof outline. One can refer to [42, Lemma 12] to understand some step(s) omitted here. We first note that the one-to-one correspondence between a central density of zero and a total mass of zero is trivial, as both correspond to the vacuum (zero function). Hence we can now turn to the case where mass is positive.
The idea of Lieb and Yau is: let $G=(0,\infty)$, and
[…]
For each $m\in G$, we know by Theorem 2.6 and Proposition 2.42 there is a unique minimizer $\rho_{m}$, and a unique central density $\alpha_{m}$ (see also Corollary 2.43). We let $\Gamma:G\rightarrow D$ denote this map from $m\in G$ to $\alpha_{m}\in D$.
Then in [42, Lemma 12], one can show $\Gamma$ (or $\Gamma^{-1}$) is a homeomorphism between $G$ and $D$, and $D$ turns out to be closed in $\mathbb{R}^{+}$. $G$ is connected, so is $D$, hence $D$ is a closed interval in $\mathbb{R}^{+}$. Since $\Gamma^{-1}$ is a homeomorphism, then $D$ can only be $\mathbb{R}^{+}$. ∎
@@ 3 Scaling Relations between Stars with Different Mass
The goal of this section is to provide a description of scaling method and scaling relations between solutions with different mass, which can help us to find more quantitative properties. We have been using those results in the previous arguments and will continue to use them in the following. We assume the polytropic equation of state $P(\rho)=K\rho^{\gamma}$ with index $\gamma>\frac{4}{3}$ holds true from this section.
@@ 3.1 Scaling relations between solutions to partial differential equations
where $\pi$ is the pressure and $v>0$ is the viscosity coefficient. For simplicity we assume $v=1$ is constant. Formally, we take the divergence of the first equation and use the second equation to eliminate some terms and obtain $\Delta\pi=-\operatorname{div}((u\cdot\nabla)u)$, i.e.
$\operatorname{div}((u\cdot\nabla)u)+\operatorname{div}(\nabla\pi)=0$
Then we can introduce Leray projection operator $\mathbb{P}=Id+\nabla(-\Delta)^{-1}\operatorname{div}$ [18] to get rid of the pressure term from the first equation and obtain:
@@
However, when we consider the reduced Euler-Poisson equations (EP’) with $\omega=0$, or the corresponding Euler-Lagrange equations (EL) with $J=0$ (non-rotating case), since the pressure $P(\rho)=K\rho^{\gamma}$ or the function $A^{\prime}(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ is not linear and there is an extra term $V_{\rho},c$ may not be chosen arbitrarily. With an abuse of notation, we assume $g_{A,B}(x)=Ag(Bx)$. Let a function $\sigma(x)$ be a solution to (EL), i.e., $A^{\prime}(\sigma(x))=\left[V_{\sigma}(x)+\lambda\right]_{+}$, then we have
$A^{\prime}\left(\sigma_{\widetilde{A},B}(x)\right)=\frac{K\gamma}{\gamma-1}\left(\sigma_{\widetilde{A},B}(x)\right)^{\gamma-1}=\widetilde{A}^{\gamma-1}A^{\prime}(\sigma(Bx))$
$V_{\sigma_{A,B}}(x)=\int_{\mathbb{R}^{3}}\frac{\widetilde{A}\sigma(By)}{|x-y|}\,dy=\widetilde{A}B^{-2}\int_{\mathbb{R}^{3}}\frac{\sigma(By)}{|Bx-By|}\,d(By)=\widetilde{A}B^{-2}V_{\sigma}(Bx)$
Hence if $\widetilde{A}^{\gamma-1}=\widetilde{A}B^{-2}$ (i.e., $B=\widetilde{A}^{\frac{2-\gamma}{2}}$), we will have $\sigma_{\widetilde{A},B}$ is also a solution to (EL):
$A^{\prime}\left(\sigma_{\widetilde{A},B}(x)\right)=\left[V_{\sigma_{\bar{A},B}}(x)+\lambda_{\widetilde{A},B}\right]_{+}$
with $\lambda_{\widetilde{A},B}=\widetilde{A}B^{-2}$.
Moreover, if we know $g$ has mass $m$ (again without loss of generality we assume $m\neq 0$ unless otherwise specified) and hope $g_{A,B}$ has mass 1. Then $\int_{\mathbb{R}^{3}}g_{A,B}\,dx=AB^{-3}$ $\int_{\mathbb{R}^{3}}g(Bx)\,d(Bx)=AB^{-3}m=1$. Therefore, we can solve $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$.
@@
For the case of Euler-Lagrange equations (EL), one can see the arguments above. For the case of the reduced Euler-Poisson equations (EP’), notice (EP’) becomes
$(\nabla P(g(x)))-g(x)\left(\nabla V_{g}(x)\right)=0$
We also notice
$\left(\nabla P\left(g_{A,B}(x)\right)\right)-g_{A,B}(x)\left(\nabla V_{g_{A,B}}(x)\right)=A^{\gamma}B\left(\nabla_{Bx}P(g(Bx))\right)-A^{2}B^{-1}g(Bx)\left(\nabla_{Bx}V_{g}(Bx)\right)$
thus the arguments are essentially the same since we have ${A}^{\gamma-2}B^{2}=1$. ∎
Therefore, to solve (EL), we can also assume a solution has mass 1. Notice the radial solution with certain mass is actually a minimizer and unique (see Proposition 2.42 and Proposition 2.46) and vice versa. It can help us to discuss more properties and relations between minimizers with different mass in the next subsection.
@@ 3.2 Scaling relations between minimizers of $E_{0}(\rho)$ with different mass
Recall (5) that $E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}=\int_{\mathbb{R}^{3}}A(\rho)\,dx-\frac{1}{2}\iint_{{\mathbb{R}^{3}}\times{\mathbb{R}^{3}}}\frac{\rho(x)\rho(y)}{|x-y|}\,dxdy$, where $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$.
Let $\sigma$ be in the admissible class (4), then we have
$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int_{\mathbb{R}^{3}}A\left(\sigma_{\widetilde{A},B}\right)\,dx$
$=\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}\,dx$
$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}\,d(Bx)$
@@
$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint_{{\mathbb{R}^{3}}\times{\mathbb{R}^{3}}}\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}\,dxdy$
$=\widetilde{A}^{2}B^{-5}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}\,d(Bx)\,d(By)$
@@
$\int_{\mathbb{R}^{3}}\sigma_{\widetilde{A},B}\,dx$ $=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma(Bx)\,d(Bx)$
$=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma\,dx$
Assume $\sigma_{m}$ is a minimizer of $E_{0}(\rho)$ on $m{R}\left(\mathbb{R}^{3}\right)$, if $A^{\gamma}B^{-3}=A^{2}B^{-5}$ (i.e., $B=A^{\frac{2-\gamma}{2}}$), we will have $\left(\sigma_{m}\right)_{A,B}$ is a minimizer of $A^{\gamma}B^{-3}E_{0}(\rho)$ on $AB^{-3}m{R}\left(\mathbb{R}^{3}\right)$ (one can show it by contradiction arguments), and thereby a minimizer of $E_{0}(\rho)$ on $AB^{-3}m{R}\left(\mathbb{R}^{3}\right)$. If we hope $\left(\sigma_{m}\right)_{A,B}$ has mass $1$, then $AB^{-3}m=1$. Therefore, we can solve $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, which are compatible with the results in subsection 3.1 as one can expect.
@@
Thanks to Theorem 2.6, we know $\|\sigma\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ and spt $\sigma$ is contained in a ball of radius $R_{1}$, therefore, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{m}=\frac{C_{1}}{A}$ and spt $\sigma_{m}$ is contained in a ball of radius $R_{m}=BR_{1}$. Since $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, if we further assume $\gamma>2$, we know $\lim\limits_{m\rightarrow 0}A=+\infty$ and $\lim\limits_{m\rightarrow 0}B=0$, thus $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ and the size of $\sigma_{m}$’s support will go to 0 when $m\rightarrow 0$, with rates $m^{\frac{2}{3\gamma-4}}$ and $m^{\frac{\gamma-2}{3\gamma-4}}$ respectively. Notice $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus it gives the decay rate and strengthens Corollary 2.43. However, when $\gamma<2$, $\lim\limits_{m\rightarrow 0}B=\infty$, the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread). It coincides with a result mentioned in Lieb and Yau’s paper [42, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$. In their paper quantum mechanics (fermions case) is discussed.
@@
Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
where $\rho(x)\coloneq\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$, $\rho$ has mass 1.
@@
The result $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$-differentiable at $\rho_{m}$ comes from Lemma 2.9. Also easy to check $U(\rho)<\infty$ and has mass 1, and $E_{0}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Given $\tau_{m}\in P_{\infty}\left(\rho_{m}\right)$, one can check $\tau:=\left(\tau_{m}\right)_{A,B}$ is in $P_{\infty}(\rho)$. By the definition of variational derivatives, one has
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\rho_{m}\right)\tau_{m}\,dx$ $=\lim_{t\rightarrow\infty}\frac{E_{0}\left(\rho_{m}+t\tau_{m}\right)-E_{0}\left(\rho_{m}\right)}{t}$
@@
$=A^{-\gamma}B^{3}\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)\tau\,dx$
$=A^{-\gamma-1}\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)(x)\tau_{m}(Bx)\,d(Bx)$
$=A^{-\gamma-1}\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)\tau_{m}(x)\,dx$
Since there are enough functions $\tau_{m}$ in $P_{\infty}\left(\rho_{m}\right)$, thus we have
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
@@
Suppose $U(\rho_{m})<\infty$, there is another way to show the result. Let $\rho(x)=A\rho_{m}(Bx)$ with $A=m^{-\frac{2}{3\gamma-4}}$ and $B=m^{\frac{\gamma-2}{3\gamma-4}}$, we already know $E_{0}^{\prime}(\rho)=A^{\prime}(\rho(x))-V_{\rho}(x)$, and then one can also compute explicitly that
$A^{\prime}\left(\rho_{m}\right)(x)=A^{\prime}\left(\frac{1}{A}\rho\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}A^{\prime}(\rho)\left(\frac{1}{B}x\right)$
$V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\rho}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\rho}\left(\frac{1}{B}x\right)$
and obtain again
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=\frac{1}{A^{\gamma-1}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
Appendix
Appendix A Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this paper.
Proposition A.1 (Hardy-Littlewood-Sobolev Inequality [5, Theorem 1.7]).